# footnotetext: last modified on March 7, 2024footnotetext: MSC classification: 46L35, 20F67 footnotetext: Keywords: twisted group C*-algebra, acylindrically hyperbolic group, stable rank one

### 07 Mar 2024 v1

@@ Acknowledgements
The author thanks Hannes Thiel for inspiring conversations about the stable rank of group C*-algebras, motivating us to write down the present note. We also thank Dennis Osin for pointing out improvements to the introduction of an earlier version of this work.
@@ 2 Proofs
Let $K\unlhd\Gamma$ be the finite radical and write $\Lambda=\Gamma/K$. Then by [MO15, Lemma 3.9] (see also [MO19, Lemma 1]) it follows that $\Lambda$ is acylindrically hyperbolic with trivial finite radical. After choosing a section $\mathrm{s}\colon\Lambda\to\Gamma$ satisfying $\mathrm{s}(e)=e$, we infer from [PR89, Theorem 4.1] that there is an identification with a twisted crossed product $\mathrm{C}^{*}_{\mathrm{red}}(\Gamma,\sigma)\cong\mathbb{C}[K,\sigma]\rtimes_{\alpha,\rho,\mathrm{red}}\Lambda$. Here the maps $\alpha\colon\Lambda\to\mathrm{Aut}(\mathbb{C}[K,\sigma])$ and $\rho\colon\Lambda\times\Lambda\to\mathcal{U}(\mathbb{C}[K,\sigma])$ are given on $h,h_{1},h_{2}\in\Lambda$ and $k\in K$ by the formulas

### 11 Mar 2024 v2

# footnotetext: last modified on March 11, 2024footnotetext: MSC classification: 46L35, 20F67 footnotetext: Keywords: twisted group C*-algebra, acylindrically hyperbolic group, stable rank one

@@ 1 Introduction
Upon publication of this note, the author was informed that Bédos-Omland had made a similar observation in [BO18, Appendix A, proof of Lemma A.2], inferring stable rank one for certain twisted group C*-algebras. This result could alternatively be used in our proof of Theorem 1. However, as the formulation of Proposition 2 also leads to the observation made in Corollary 5, we decided to keep it in its present form.
@@ Acknowledgements
The author thanks Hannes Thiel for inspiring conversations about the stable rank of group C*-algebras, motivating us to write down the present note. We also thank Dennis Osin for pointing out improvements to the introduction of an earlier version of this work. We thank Jamie Bell and Tron Omland for pointing out the Appendix of [BO18] and Tron Omland for pointing out [Béd91].
@@ 2 Proofs
Let $K\unlhd\Gamma$ be the finite radical and write $\Lambda=\Gamma/K$. Then by [MO15, Lemma 3.9] (see also [MO19, Lemma 1]) it follows that $\Lambda$ is acylindrically hyperbolic with trivial finite radical. After choosing a section $\mathrm{s}\colon\Lambda\to\Gamma$ satisfying $\mathrm{s}(e)=e$, we infer from [Béd91, Theorem 2.1] (see also [PR89, Theorem 4.1]) that there is an identification with a twisted crossed product $\mathrm{C}^{*}_{\mathrm{red}}(\Gamma,\sigma)\cong\mathbb{C}[K,\sigma]\rtimes_{\alpha,\rho,\mathrm{red}}\Lambda$. Here the maps $\alpha\colon\Lambda\to\mathrm{Aut}(\mathbb{C}[K,\sigma])$ and $\rho\colon\Lambda\times\Lambda\to\mathcal{U}(\mathbb{C}[K,\sigma])$ are given on $h,h_{1},h_{2}\in\Lambda$ and $k\in K$ by the formulas